# Orthogonal projectors of binary LCD codes

### v2

@@ Abstract
We prove that binary even LCD code and some graphs are in one-to-one correspondence in a certain way. Furthermore, we show that adjacency matrices of non-isomorphic simple graphs give inequivalent binary LCD codes, and vice versa.
@@ 1 Introduction
The interplay between codes and designs have provided many useful results [Assmus1992-rb]. Concerning relations between codes and graphs, several results are known: The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code. Since any orthogonal projector is symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors. Based on this observation, Rodrigues and Keys studied a relation between LCD codes and adjacency matrices of graphs. They gave two conditions where adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
In this paper, we consider binary LCD codes in particular, and study a relation between binary LCD codes and adjacency matrices of simple graphs. It is known that the orthogonal projector $A$ of a binary LCD code is a symmetric matrix with $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Hence any adjacency matrix satisfies the latter condition is an orthogonal projector of binary LCD code, as shown in [Key2018-ec, Brouwer1992-nc]. We show that such adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic. This establishes a one-to-one correspondence between binary even LCD codes and the above-mentioned class of graphs. Haemers, Peeters and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs. For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes. This paper gives an explanation to the above observation. Furthermore, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD $[n,k,d]$ codes with the following parameters:
$\begin{split}(n,k,d)=&(31,10,10),(31,11,10),(31,20,6),(33,10,12),\\ &(33,11,11),(33,12,10),(33,13,10),(33,20,6),\\ &(33,22,6),(33,21,6),(34,13,10),(34,22,6).\end{split}$
We note that orthogonal projectors of the above optimal codes are adjacency matrices of regular graphs. By the above example, we show that from optimal binary even LCD codes we can generate graphs whose adjacency matrices give codes with large minimum weight.
This paper is organized as follows: In Section 2, we introduce definitions and basic results of LCD codes and strongly regular graphs. In Section 3, we prove the main results. In Section 4, we apply the main results to a study of linear codes obtained as row spans of adjacency matrices of strongly regular graphs, and to a construction of optimal binary LCD codes.
@@ 2 Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of order $q$, where $q$ is a prime power. Let $\mathbb{F}_{q}^{n}$ be the vector space of all $n$-tuples over $\mathbb{F}_{q}$. A $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ is said to be an $[n,k]$ code over $\mathbb{F}_{q}$. Especially, codes over $\mathbb{F}_{2}$ are said to be binary codes. Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. The parameters $n$ and $k$ are said to be the length and the dimension of $C$, respectively. A vector in $C$ is called a codeword. The weight of $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{F}_{q}^{n}$ is defined as $\operatorname{wt}(x)=|\{i\mid x_{i}\neq 0\}|$. The minimum weight $d(C)$ of $C$ is the minimum weight among all nonzero codewords in $C$. If the minimum weight of $C$ equals to $d$, then $C$ is said to be an $[n,k,d]$ code over $\mathbb{F}_{q}$. We say that a code $C$ is even if all codewords have even weights; otherwise, the code $C$ is odd. Two $[n,k]$ codes $C_{1}$ and $C_{2}$ over $\mathbb{F}_{q}$ are equivalent if there is a monomial matrix $M$ such that $C_{2}=\{cM\mid c\in C_{1}\}$. The equivalence of two codes $C_{1}$ and $C_{2}$ is denoted by $C_{1}\simeq C_{2}$. A generator matrix of a code $C$ is any matrix whose rows form a basis of $C$.
@@
Let $C$ be a code over $\mathbb{F}_{q}$ with length $n$ and let $T\subseteq\{1,2,\dots,n\}$. Deleting every coordinate $i\in T$ in every codeword of $C$ gives a code called the punctured code on $T$ and denoted by $C^{T}$. Define a subcode $C(T)=\{(c_{1},c_{2},\dots,c_{n})\in C\mid c_{i}=0\text{ for all }i\in T\}$. Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$. If $|T|=1$, say $T=\{i\}$, then we will write $C^{\{i\}}$ and $C_{\{i\}}$ simply as $C^{i}$ and $C_{i}$, respectively. Bouyuklieva gave a characterization of both punctured codes and shortened codes of binary LCD codes:
@@
A graph $\Gamma=(V,E)$ consists of a finite set of vertices $V$ together with a set of edges $E\subset\binom{V}{2}$. In this paper, we are concerned only with simple graphs, namely, graphs which are undirected, and without loops or multiple edges. We say that $x,y\in V$ are adjacent if $\{x,y\}\in E$, and denote by $x\sim y$. A vertex $y\in V$ is a neighbor of $x\in V$ if $x$ and $y$ are adjacent. The cardinality of the set of neighbors $N(x)=\{y\in V\mid x\sim y\}$ is called the valency of $x\in V$. A graph is $k$-regular if all its vertices have the same valency $k$. A simple regular graph is strongly regular with parameters $(v,k,\lambda,\mu)$ if it has $v$ vertices, valency $k$, and if any two adjacent vertices are together adjacent to $\lambda$ vertices, while any two nonadjacent vertices are together adjacent to $\mu$ vertices. A strongly regular graph with parameters $(v,k,\lambda,\mu)$ is denoted by $srg(v,k,\lambda,\mu)$. An adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labelled by the vertices $x,y\in V$, defined by $A_{x,y}=1$ if $x\sim y$ and $A_{x,y}=0$ otherwise. Two graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ are equivalent if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
@@
The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of some strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained several computational results for strongly regular graphs with small vertices. Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors. Especially, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
@@ 3 Main results
In this section, we give main results on a correspondence between binary even LCD codes and adjacency matrices of some simple graphs. Also, we provide a lower bound on the minimum weights of binary codes that come from strongly regular graphs.
3.1 Characterization
@@
Let $C$ be a binary LCD code with orthogonal projector $A$. Then $C$ is the row span of $A$.
@@
Let $C$ and $C^{\prime}$ be binary LCD codes with orthogonal projectors $A$ and $A^{\prime}$, respectively. Then, $C$ is equivalent to $C^{\prime}$ if and only if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$.
@@
Let $G$ and $G^{\prime}$ be generator matrices of $C$ and $C^{\prime}$ respectively. If $C$ and $C^{\prime}$ are equivalent, then there are a monomial matrix $M$ and a permutation matrix $P$ such that $G^{\prime}=MGP$. By Theorem 2.1, $A^{\prime}=P^{T}AP$ follows.
Conversely, suppose that there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$. Any permutation of the rows of $A$, $P^{T}$ in particular, preserves its row span. By Lemma 3.1, the row spans of $A$ and $A^{\prime}$ are $C$ and $C^{\prime}$ respectively. It follows that $C^{\prime}=CP$, and hence $C$ and $C^{\prime}$ are equivalent, as required. ∎
Corollary 3.3.
Let $C$ and $C^{\prime}$ be binary LCD codes with orthogonal projectors $A$ and $A^{\prime}$, respectively. If the traces of $A$ and $A^{\prime}$ are distinct, then $C$ and $C^{\prime}$ are inequivalent.
Remark 3.4.
By the proof of Lemma 3.2, it follows that the orthogonal projector of an LCD code is independent of the choice of a generator matrix.
Remark 3.5.
@@
Lemma 3.6.
Let $C$ be a binary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of a simple graph if and only if $C$ is an even code.
Proof.
Let $r_{i}$ be the $i$-th row of $A$, and let $n$ be the length of $C$. If $A$ is an adjacency matrix of a simple graph, then any diagonal entry of $A$ is zero. This fact, along with $A^{2}=A$, implies $(r_{i},r_{i})=0$ for all $1\leq i\leq n$. That is, any row of $A$ has an even weight. By Lemma 3.1, $C$ is even.
Conversely, suppose $C$ is even. Since the orthogonal projector $A$ is a symmetric $(0,1)$-matrix, it remains to show that any diagonal entry of $A$ is 00. This can be done by the above argument along with the fact that every codeword of $C$ has an even weight and $A^{2}=A$. This completes the proof. ∎
Theorem 3.7.
Proof.
This immediately follows from Lemmas 3.1, 3.2 and 3.6. ∎
@@
3.2 Bounds on minimum weights
In this subsection, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs. These results will be applied in Section 4. Note that all the assertions in this subsection hold not only for binary LCD codes but also for any binary codes.

### v3

# Orthogonal projectors of LCD codes and their graph representations

@@ Abstract
We establish one-to-one correspondences between (i) binary even LCD codes and certain simple graphs, and (ii) ternary LCD codes and certain two-graphs.
@@ 1 Introduction
The interplay between codes and designs has provided many useful results [Assmus1992-rb]. Concerning relations between codes and graphs, several results are known: The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code. Since any orthogonal projector is a symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors. Based on this observation, Rodrigues and Keys studied the relation between LCD codes and adjacency matrices of graphs. They gave two conditions under which adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
In this paper, we establish one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs. It is known that the orthogonal projector $A$ of a binary LCD code is a symmetric matrix with $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Hence any adjacency matrix satisfies the latter condition is an orthogonal projector of binary LCD code, as shown in [Key2018-ec, Brouwer1992-nc]. We show that such an adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic. Furthermore, we prove an analogous results for the relation between ternary LCD codes and two-graphs. Haemers, Peeters, and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs. For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes. This paper gives an explanation to the above observation.
The remaining part of this paper is organized as follows: In Section 2, we introduce definitions and basic results of LCD codes, simple graphs, and two-graphs. We give our main results in Section 3. Section 4 establishes auxiliary results on the minimum weight of codes derived from strongly regular graphs, which are later used in Section 5. In that section, we apply our main results to analyze binary LCD codes obtained as row spans of adjacency matrices of strongly regular graphs. Finally, Section 6 presents concluding remarks and open problems.
@@ 2 Preliminaries
2.1 LCD codes
Let $\mathbb{F}_{q}$ be the finite field of order $q$, where $q$ is a prime power. Let $\mathbb{F}_{q}^{n}$ be the vector space of all $n$-tuples over $\mathbb{F}_{q}$. A $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ is said to be an $[n,k]$ code over $\mathbb{F}_{q}$. In particular, codes over $\mathbb{F}_{2}$ and $\mathbb{F}_{3}$ are said to be binary and ternary codes, respectively. Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. The parameters $n$ and $k$ are said to be the length and the dimension of $C$, respectively. A vector in $C$ is called a codeword. The weight of $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{F}_{q}^{n}$ is defined as $\operatorname{wt}(x)=|\{i\mid x_{i}\neq 0\}|$. The minimum weight $d(C)$ of $C$ is the minimum weight among all nonzero codewords in $C$. If the minimum weight of $C$ equals to $d$, then $C$ is said to be an $[n,k,d]$ code over $\mathbb{F}_{q}$. We say that a code $C$ is even if all codewords have even weights; otherwise, the code $C$ is odd. Two $[n,k]$ codes $C_{1}$ and $C_{2}$ over $\mathbb{F}_{q}$ are equivalent if there is a monomial matrix $M$ such that $C_{2}=\{cM\mid c\in C_{1}\}$. The equivalence of two codes $C_{1}$ and $C_{2}$ is denoted by $C_{1}\simeq C_{2}$. A generator matrix of a code $C$ is a matrix whose rows form a basis of $C$.
The orthogonal projector $\Pi_{C}$ of an $[n,k]$ code $C$ over $\mathbb{F}_{q}$ is a unique linear map from $\mathbb{F}_{q}^{n}$ to $C$ satisfying
$v\Pi_{C}=\begin{cases}v&\textrm{if }v\in C,\\ 0&\textrm{otherwise}.\end{cases}$
An orthogonal projector is represented by a matrix $A$ satisfying $A^{T}=A$ and $A^{2}=A$. Conversely, any matrix satisfying $A^{T}=A$ and $A^{2}=A$ represents the orthogonal projector of some code. We refer the reader to [massey_linear_1992, bardet2019permutation] for more information on orthogonal projectors. In this paper, we often identify the orthogonal projector with its matrix representation in the standard basis.
@@ 2.1 LCD codes
Let $C$ be a code over $\mathbb{F}_{q}$ with length $n$ and let $T\subseteq\{1,2,\dots,n\}$. Deleting every coordinate $i\in T$ in every codeword of $C$ gives a code called the punctured code on $T$ and denoted by $C^{T}$. Define a subcode $C(T)=\{(c_{1},c_{2},\dots,c_{n})\in C\mid c_{i}=0\text{ for all }i\in T\}$. Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$. We will write $C^{\{i\}}$ and $C_{\{i\}}$ simply as $C^{i}$ and $C_{i}$, respectively. Bouyuklieva [bouyuklieva_optimal_2021] gave a characterization of both punctured codes and shortened codes of binary LCD codes:
@@ 2.2 Strongly regular graphs
In this subsection, we review the definition and basic properties of strongly regular graphs. For a comprehensive treatment on this topic, we refer the reader to [cameron1991designs]. A graph $\Gamma=(V,E)$ consists of a finite set of vertices $V$ together with a set of edges $E\subset\binom{V}{2}$. In this paper, we are concerned only with simple graphs, namely, graphs which are undirected, and without loops or multiple edges. We say that $x,y\in V$ are adjacent if $\{x,y\}\in E$, and denote by $x\sim y$. A vertex $y\in V$ is a neighbor of $x\in V$ if $x$ and $y$ are adjacent. The cardinality of the set of neighbors $N(x)=\{y\in V\mid x\sim y\}$ is called the valency of $x\in V$. A graph is $k$-regular if all its vertices have the same valency $k$. A simple regular graph is strongly regular with parameters $(v,k,\lambda,\mu)$ if it has $v$ vertices, valency $k$, and if any two adjacent vertices are together adjacent to $\lambda$ vertices, while any two nonadjacent vertices are together adjacent to $\mu$ vertices. A strongly regular graph with parameters $(v,k,\lambda,\mu)$ is denoted by $srg(v,k,\lambda,\mu)$. An adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labeled by the vertices $x,y\in V$, defined by $A_{x,y}=1$ if $x\sim y$ and $A_{x,y}=0$ otherwise. Two graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ are isomorphic if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$. The elementary counting argument gives the following theorem (see [cameron1991designs]).
Theorem 2.4.
Let $A$ be an adjacency matrix of $srg(v,k,\lambda,\mu)$. Then $A^{2}=kI+\lambda A+\mu(J-I-A)$, where $I$ an $J$ denote the identity matrix and the all-ones matrix of suitable order, respectively.
@@
The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of some strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained several computational results for strongly regular graphs with small vertices. Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors. In particular, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
2.3 Two-graphs
In this subsection, we review the definition and basic properties of two-graphs. For a comprehensive treatment on this topic, we refer the reader to [cameron1991designs]. A two-graph $(V,\Delta)$ consists of a set $V$ whose elements are called vertices, and a collection $\Delta$ of unordered triples of vertices such that each $4$-tuple of $V$ contains an even number of elements of $\Delta$ as subsets. Given a simple graph $\Gamma=(V,E)$, the set of triples of $V$ whose induced subgraph in $\Gamma$ contains an odd number of edges gives rise to a two-graph $(V,\Delta)$, and every two-graph can be represented in this way (see [cameron1991designs]).
Let us introduce another adjacency matrix of a simple graph called “$\mp 1$ adjacency matrix”: A $\mp 1$ adjacency matrix $A$ of a graph $\Gamma=(V,E)$ is a $|V|\times|V|$ matrix with rows and columns labeled by the vertices $x,y\in V$, defined by $A_{x,y}=-1$ if $x\sim y$ and $A_{x,y}=1$ otherwise. It is known that simple graphs $\Gamma$ and $\Gamma^{\prime}$, with $\mp 1$ adjacency matrices $A$ and $A^{\prime}$ respectively, represent the same two-graph if there is a diagonal $(1,-1)$-matrix $D$ such that $A^{\prime}=DAD$. This defines an equivalence class of graphs, called the switching class, and hence two-graphs can be identified with the switching class of simple graphs. The two-graphs $(V,\Delta)$ and $(V^{\prime},\Delta^{\prime})$ are isomorphic if there is a bijection $V\to V^{\prime}$ that induces a bijection $\Delta\to\Delta^{\prime}$. Let $\Gamma$ and $\Gamma^{\prime}$ be two simple graphs with $\mp 1$ adjacency matrices $A$ and $A^{\prime}$, respectively. By definition, these simple graphs represent isomorphic two-graphs if there are a permutation matrix $P$ and a diagonal $(1,-1)$-matrix such that $A^{\prime}=P^{T}DADP$.
@@ 3 Main results
In this section, we establish one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs. In what follows, we prove general results on orthogonal projectors of LCD codes over $\mathbb{F}_{q}$. Specific results on binary and ternary LCD codes will be stated in the following subsections.
@@
Let $C$ be an LCD code over $\mathbb{F}_{q}$ with orthogonal projector $A$. Then $C$ equals the row span of $A$.
@@
Theorem 3.2.
Let $C$ and $C^{\prime}$ be LCD codes over $\mathbb{F}_{q}$ with orthogonal projectors $A$ and $A^{\prime}$, respectively. Then, $C$ is equivalent to $C^{\prime}$ if and only if there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$.
Proof.
Let $G$ and $G^{\prime}$ be generator matrices of $C$ and $C^{\prime}$ respectively. If $C$ and $C^{\prime}$ are equivalent, then there are monomial matrices $M$ and $N$ such that $G^{\prime}=MGN$. Now the following computation shows that $A^{\prime}=N^{T}AN$.
$\begin{split}A^{\prime}&=(MGN)^{T}((MGN)(MGN)^{T})^{-1}(MGN)\\ &=N^{T}G^{T}M^{T}(MGG^{T}M^{T})^{-1}(MGN)\\ &=N^{T}G^{T}(GG^{T})^{-1}GN\\ &=N^{T}AN.\\ \end{split}$
Conversely, suppose that there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$. The row span of $A$ is preserved by both permuting coordinates and multiplying any coordinate by nonzero scalar. Thus multiplying $A$ by $N^{T}$ from left preserves the row span of $A$. Lemma 3.1 tells us that the row spans of $A$ and $A^{\prime}$ are $C$ and $C^{\prime}$, respectively. It follows that $C^{\prime}=CN$, and hence $C$ and $C^{\prime}$ are equivalent, as required. ∎
Remark 3.3.
An analogous statement about permutation equivalence was presented in [bardet2019permutation] in the context of public-key cryptography.
@@
Let $C$ be an LCD code of length $n$ over $\mathbb{F}_{q}$ with orthogonal projector $A$, and let $r_{i}$ denote the $i$-th row of $A$. Then every diagonal entry of $A$ is zero if and only if $(r_{i},r_{i})=0$ for all $1\leq i\leq n$.
@@
This is straightforward for the $i$-th diagonal entry of $A^{2}$ equals $(r_{i},r_{i})$. ∎
3.1 Binary LCD codes and simple graphs
In this subsection, we restrict ourselves to binary codes. Our aim is to establish the equivalence between binary even LCD codes and certain simple graphs.
Lemma 3.5.
Let $C$ be a binary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of some simple graph if and only if $C$ is an even code.
Proof.
If $A$ is an adjacency matrix of a simple graph, then any diagonal entry of $A$ is zero. By Lemma 3.4, this implies that any row of $A$ has even weight. Now, the result follows by Lemma 3.1.
Conversely, suppose $C$ is even. Since the orthogonal projector $A$ is a symmetric $(0,1)$-matrix, it remains to show that any diagonal entry of $A$ is 00. Since $C$ is even, any codeword of $C$ has even weight. In particular, every row of $A$ has even weight, and the result follows by Lemma 3.4. This completes the proof. ∎
Theorem 3.6.
Proof.
The first part of the theorem is an immediate consequence of Lemmas 3.1 and 3.5. To prove the second part, recall that simple graphs $\Gamma$ and $\Gamma^{\prime}$ with adjacency matrices $A$ and $A^{\prime}$ respectively are isomorphic if there is a permutation matrix $P$ such that $A^{\prime}=P^{T}AP$. Since any permutation matrix is a monomial matrix, Theorem 3.2 implies the result. ∎
@@ 3.1 Binary LCD codes and simple graphs
3.2 Ternary LCD codes and two-graphs
In this subsection, we restrict ourselves to ternary codes. Our aim is to establish the equivalence between ternary LCD codes and certain two-graphs. Throughout this subsection, the term “adjacency matrix” means the $\mp 1$ adjacency matrix (see Section 2).
Lemma 3.8.
Let $C$ be a ternary LCD code with orthogonal projector $A$. Then $A$ is an adjacency matrix of some two-graph if and only if every row of $A$ has weight divisible by three.
Proof.
If $A$ is an adjacency matrix of a two-graph, then any diagonal entry of $A$ is zero. The result follows by Lemma 3.4 and the fact that $(x,x)=0$ if and only if $\operatorname{wt}(x)\equiv 0\pmod{3}$ for any $x\in\mathbb{F}_{3}^{n}$.
Conversely, suppose that any row of $A$ has weight divisible by three. Since the orthogonal projector $A$ is a symmetric matrix whose entry is in $\{0,1,-1\}$, it remains to show that any diagonal entry of $A$ is 00. Again, the result follows by Lemma 3.4 and the fact that $(x,x)=0$ if and only if $\operatorname{wt}(x)\equiv 0\pmod{3}$ for any $x\in\mathbb{F}_{3}^{n}$. This completes the proof. ∎
In both binary and ternary cases, the conditions for orthogonal projectors to be adjacency matrices are characterized by weights of their row vectors. In binary case, Lemma 3.1 implies every codeword to have even weight, while an analogous statement is false in ternary case. In fact, if every codeword in $C$ has weight divisible by three, then $C$ is self-dual code and hence cannot be LCD. The condition “every row of the orthogonal projector has weight divisible by three” is not insightful since such a statement is seemingly not related to any property of the corresponding code. From this perspective, the following sufficient condition is more useful.
Theorem 3.9.
Let $C$ be a ternary LCD code of length $n$ with orthogonal projector $A$. If $A$ is an adjacency matrix of some two-graph, then the dimension of $C$ is divisible by three.
Proof.
In this proof, we regard $A$ as a matrix over $\mathbb{F}_{3}$. Since $A^{2}=A$, the eigenvalue of $A$ is either 00 or $1$. To see this, let $a$ be any eigenvalue of $A$ and $v$ be an eigenvector with respect to $a$. Then $0=(A^{2}-A)v=(a^{2}-a)v$, and hence $a\in\{0,1\}$, as required. Therefore, there is an invertible matrix $P$ over $\mathbb{F}_{3}$ such that $P^{-1}AP=\operatorname{diag}{(1,\dots,1,0,\dots,0)}$. Let $D$ denote this diagonal matrix. The code $C$ is exactly the $A$-invariant subspace, so its dimension equals the number of $1$s in $D$, which we denote by $k$. Now, the proof is concluded by the facts $\operatorname{tr}(A)=\operatorname{tr}(D)$ and $\operatorname{tr}(D)\equiv k\pmod{3}$, where in the latter we regard $\operatorname{tr}(D)$ as an integer. ∎
The converse of the above theorem is false: Let $C$ be a ternary code whose generator matrix is of the form $G=\begin{pmatrix}I_{k}&A\end{pmatrix},$ where $k\equiv 0\pmod{3}$ and $A$ is $k\times(n-k)$ matrix with $AA^{T}=O$. Since $GG^{T}=I_{k}$ is invertible, $C$ is LCD, and one can easily check that
$G^{T}(GG^{T})^{-1}G=\begin{pmatrix}I_{k}&A\\ A^{T}&O\end{pmatrix}.$
Because of the k $1$s in its diagonal, this matrix fails to be an adjacency matrix of any two-graph.
Theorem 3.10.
The orthogonal projector $A$ of a ternary LCD code is an adjacency matrix of a simple graph satisfying $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Conversely, such an adjacency matrix is the orthogonal projector of a ternary LCD code. Furthermore, two binary even LCD codes are equivalent if and only if corresponding graphs are isomorphic.
Proof.
The first part of the theorem is an immediate consequence of Lemmas 3.1 and 3.5. To prove the second part, let $\Gamma$ and $\Gamma^{\prime}$ be two-graphs with adjacency matrices $A$ and $A^{\prime}$, and suppose their rows generate ternary LCD codes $C$ and $C^{\prime}$, respectively. If $\Gamma$ and $\Gamma^{\prime}$ are isomorphic, then there are a diagonal $(1,-1)$-matrix $D$ and a permutation matrix $P$ such that $A^{\prime}=P^{T}DADP$. If we write $N=DP$, then $N$ is a monomial matrix and $A^{\prime}=N^{T}AN$. Hence $C$ and $C^{\prime}$ are equivalent by Theorem 3.2. Conversely, if $C$ and $C^{\prime}$ are equivalent, then there is a monomial matrix $N$ such that $A^{\prime}=N^{T}AN$, again by Theorem 3.2. Recall that any monomial matrix can be represented by a diagonal matrix and a permutation matrix. Hence there are a diagonal $(1,-1)$-matrix $D$ and a permutation matrix $P$ such that $N=DP$. Therefore, we have $A^{\prime}=P^{T}DADP$, as required. ∎
4 Bounds on minimum weights
In this section, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs. These results will be applied in Section 5. Note that all assertions in this subsection hold not only for binary LCD codes but also for any binary codes.